OpenStudy (anonymous):

Help please! Solve the definite integral. integral of xsinx^2dx from 0 to sqrt 4x

OpenStudy (anonymous):

u = x^2

OpenStudy (anonymous):

I'm assuming that the integral is: \[\int\limits x Sin(x^2)dx\]

OpenStudy (anonymous):

so it would be integral of sin(u) du

OpenStudy (anonymous):

There should be an additional constant in there, but close :)

OpenStudy (anonymous):

ok so it would be the integral of sin which is cos and that becomes cos(u) +c and that then becomes.... cos(x^2)+c

OpenStudy (anonymous):

\[\int\limits x Sin(x^2)dx = \frac{1}{2}\int\limits \sin(u)du = \frac{1}{2}[-\cos(u)]+C\] Since you have bounds, we don't really care about the C, though.

OpenStudy (anonymous):

ok so sqrt 4pi would be first :)

OpenStudy (anonymous):

Well, your bounds were for x. You either need to find what u is when x = those things, and use those as your new bounds, or substitute back in for u.

OpenStudy (anonymous):

ok so it would be 1/2 -cos (x^2) since u is x^2

OpenStudy (anonymous):

Just to be clear, the 1/2 is multiplying the (-cos(u)). Rewritten for clarity: \[\frac{-1}{2}Cos(u)\]

OpenStudy (anonymous):

yeah sorry that is what i mean :)

OpenStudy (anonymous):

Right. and u = x^2, so you plug that in, then evaluate at the given bounds.

OpenStudy (anonymous):

so then it is sqrt of 4pi first which when squared becomes 4pi and -cos(4pi) = -1 and divided by 2 = -0.5

OpenStudy (anonymous):

subtract -cos(0)/2 from that, which equals -0.5 and the answer is 0 :)

OpenStudy (anonymous):

Looks good to me :)

OpenStudy (anonymous):

yay! :D thank you!

OpenStudy (anonymous):

My pleasure :)